OpenStudy (anonymous):

How would I simplify -√189 ?

OpenStudy (tkhunny):

You go on a factor hunt. Can you think of ANY whole number that divides 189 evenly? I suggest you start with 3.

OpenStudy (anonymous):

\[-\sqrt{189}=-\sqrt{9*21}=-\sqrt{9}*\sqrt{21}=-3\sqrt{21}\]

OpenStudy (anonymous):

@KeithAfasCalcLover is correct

OpenStudy (tkhunny):

You would be more correct if you had done the work,

OpenStudy (anonymous):

I suppose you could simplify it further... \[-3\sqrt{21}=-3\sqrt{3*7}=-3*\sqrt{3}*\sqrt{7}=-(3)^{\frac{3}{2}}\sqrt{7}\]

OpenStudy (anonymous):

Thanks a whole lot @KeithAfasCalcLover. Life saver. :3

OpenStudy (anonymous):

Hey my pleasure ;)